SITE 1071
(Proposed Site MAT-8B-3)

Hole 1071A
At 0009 hr on 23 June, the first beacon (14.0 kHz, 193 dB) was dropped on the dynamic global positioning system (dGPS) coordinates for prospectus Site MAT-8B-3. During the process of stabilizing the ship on site, it became apparent that the automatic station keeping (ASK) system was experiencing short, unstable periods. At the time, it was assumed that strong bottom or eddy currents were causing the beacons to drift off location or flutter. Subsequently, it was determined that the beacon was dropped 20.7 m southwest of the intended location; this was assumed to be an artifact of an unstable dGPS signal (±20 m). As the beacon's transmission signal angle provides a signal radius of ~15 m in 90 m water depth, the 20.7-m offset was considered unacceptable for shallow-water operations.

A second beacon (17.0 kHz, 193 dB), equipped with a bar through the anchor weight to prevent rolling, was dropped at 0302 hr; however, its signal was unstable. A third, more powerful beacon (15.0 kHz, 205 dB) was dropped at 0326 hr, but it had to be turned off because it was interfering with reception of the first beacon's signal. A fourth beacon (16.0 kHz, 198 dB) was dropped at 0720 hr; it landed ~12 m northeast of Hole 1071A, providing a dependable signal relative to previous beacons. The second and third beacons were released and retrieved. However, the ASK system was still experiencing ~1.5 min unstable intervals when beacon signals were poor, which was attributed to beacon motion in changing currents. Consequently, a 1-m high stand with an angle-iron square base and drill-string guide rings was constructed out of scrap casing to place a beacon exactly on site and prevent its fluctuation in currents. A fifth beacon (15.0 kHz, 193 dB) was placed in this stand, and dropped to the seafloor along the pipe with the bit on bottom. Although the ASK system was still experiencing occasional signal losses from the two remaining active beacons, one in the stand and one on a tether, the improvement in the ship's dGPS system stability affected by the ODP technical staff in the meantime allowed coring operations to commence.

The APC/XCB bottom-hole assembly (BHA) was run to 78 meters below rig floor (mbrf), followed by the vibration-isolated television (VIT) frame. A reference survey (video and sonar) was conducted on a 7.6-m square and revealed a clean, flat seafloor. Hole 1071A was spudded at 1300 hr on 23 June, cutting APC Cores 1071A-1H through 7H from 0 to 36.3 mbsf (76.4% recovery). The bit tagged a firm bottom at 98.0 mbrf. The distance between the rig floor and sea level was 11.1 m, thus, the water depth was 88.3 meters below sea level (mbsl). All seven cores were partial strokes with the bit advanced by recovery. Core 6H recovered liquified sand, which is assumed to be mostly suck-in. Core 7H had a shattered liner that had to be heated and pumped out, which required 2 hr.

XCB Cores 1071A-8X through 10X were cut from 36.3 to 51.9 mbsf with 7.0% recovery. The coring parameters were varied in an unsuccessful effort to optimize recovery. A Davis-Villinger Temperature Probe (DVTP) was run after Core 9X at 49.3 mbsf. At 2330 hr on 23 June, a ground fault was detected in the top drive. Further investigation indicated that the motor field windings had burned out. Coring was terminated and the bit was pulled to complete repairs, clearing the rotary at 0355 hr, and ending Hole 1071A. Replacement of the top drive motor with a spare carried on board required 25 hr.

Hole 1071B
During the top drive repairs, the ship was moved 10 m at 35š (along seismic Line 806 in the direction of proposed Site MAT-8B-2). The same APC/XCB BHA used at Hole 1071A was rerun with a 10-1/8-in. PDC bit. The bit was set at 93 mbrf, and an APC core was attempted at 0817 hr on 25 June; however, the BHA apparently moved and broke off the APC shoe and core barrel with no recovery. ASK excursion warnings, set at 4% (yellow light) and 8% (red light) of total water depth, shut down operations for 2.75 hr on 25 June.

Hole 1071B was spudded at 0945 hr on 25 June with an XCB core barrel. Although the beacons were now reasonably stable, the ASK system was still experiencing numerous excursions. Therefore, the decision was made to drill ahead to bury the drill collars, and put a more flexible drill pipe in the water column below the ship. Therefore, Hole 1071B was washed from 0 to 39.8 mbsf, and a 1.31 m wash core (1071B-1W) was recovered. Cores 1071B-2X through 3X were cut from 39.8 to 52.9 mbsf, recovering 2.69 m (20.5%).

At 1515 hr on 25 June, the newly installed top drive began smoking. A jammed or dragging air brake was suspected; therefore, the bit was pulled out to the seafloor at 1610 hr on 25 June. The brake recess was cleaned out and the top drive was returned to service after 3.25 hr.

The ship position had not changed, and the bit voluntarily reentered Hole 1071B at 1845 hr. The bit was run to bottom of the hole with very light reaming. A sepiolite mud sweep was conducted at 53 mbsf and XCB coring resumed. Cores 1071B-4X through 6X were recovered from 52.9 to 77.4 mbsf with 33.9% recovery. At 77.3 mbsf at 2200 hr on 25 June, the pipe became stuck when the bit was run to bottom for Core 7X. The pipe was worked to 100,000 lb overpull and 800 amps torque when the top drive failed. Unable to rotate or circulate, the pipe was pulled to 150,000 lb without success. Preparations were made to sever the pipe; however, a final effort to fracture the formation with 3000 psi pump pressure at 500 gpm and 30 bbl of sepiolite mud freed the pipe at 2245 hr. The bit was pulled to the ship for the top-drive repairs, and cleared the rotary at 0115 hr. Parts were scavenged from the original (now backup) top drive to replace a pinion gear that had slipped down off the draft shaft; repairs were completed at 1845 on 26 June after 17.5 hr downtime.

Hole 1071C
During the third top-drive repair, the ship was moved 6.7 m at 261š from Hole 1071B to Hole 1071C to permit time for ASK stability evaluation. Utilizing the same BHA and bit with an XCB core barrel, Hole 1071C was spudded at 2035 hr on 26 June. To get the BHA buried as quickly as possible, the hole was washed to 58.4 mbsf; wash Core 1071C-1W had no recovery.

XCB coring was initiated, but the drill pipe became stuck after Core 1071C-4X (83.3 mbsf), stalling the rotary; the pipe would not move with 60,000 lb overpull. The drill pipe was freed after setting weight down on the BHA and increasing pump pressure and circulation. XCB Cores 2X through 20X were cut from 58.4 to 219.0 mbsf with 29.1% recovery. A 20-bbl precautionary sepiolite mud sweep was circulated on every core to clean out loose coarse sand.

At 2230 hr on June 27, after Core 20X (219.0 mbsf), an ASK red light (8%) alert was sounded when a tidal surge caused a 30.5-m excursion. As a precaution, the pipe was set in the 500-ton elevators with a maximum pull of 10,000 lb. A wiper trip was made to 94 mbrf to examine the pipe located at the seafloor, which was essentially undamaged. The pipe was run back to bottom with no fill.

XCB Cores 1071C-21X through 24X were taken from 219.0 to 251.6 mbsf with 16.1% recovery. A heavy back flow of fluid and coarse sand pushed core barrel 25X back out of the pipe. Fluid flow initially reached the top drive located halfway up the derrick and deposited ~5 cm of coarse sand and shell debris on the rig floor. No hydrocarbons had been noted in the previous core or in the water flowing out of the pipe; therefore, there did not appear to be a safety risk either to equipment or to personnel. Despite the heavy back flow, the drill string seemed to remain free. A precautionary 20-bbl sepiolite mud sweep was circulated in an effort to remove any loose sand bridges, but it had no effect on flow. The initial conclusion was that the lockable float valve (LFV) had failed in the XCB assembly and that a charged aquifer had been penetrated, allowing fluid to vent up the pipe. However, subsequent chemical analyses of both the fluid and surface seawater indicated that the flowing water was in fact drilling fluid.

The bit was pulled to 74.8 mbsf, the core barrel was retrieved, and the VIT was deployed to view the hole at the seafloor. An 8-m wide, round crater with billowing water and sand was observed, but no strong flow was evident. Two 4-m-long pock marks indenting the side of the crater indicated where the drill pipe had been forced into the crater rim during previous ASK excursions. The BHA was then pulled to the ship, and the LFV was replaced with a flapper valve (FV), in which the flapper is normally closed (by a spring) and does not lock open to permit wireline logging. Hole 1071C was reentered easily and cleaned out to 251.6 mbsf through a few minor soft bridges.

XCB Cores 25X through 27X were recovered from 251.6 to 257.4 mbsf. A 0.3-m-thick well-indurated sandstone was recovered in Core 26X; the core barrel was pulled early because of the slow rate of penetration (ROP; 0.4 m/hr). After cutting Core 27X through an estimated 0.25 m of the sandstone and 2-3 m of very soft, presumably sandy, material, the pipe became stuck at 257.4 mbsf at 2115 hr on 28 June. The pipe was moved up 2 m with overpulls of up to 220,000 lb in an unsuccessful attempt to free the drill string. Wireline logging through the pipe was considered, but logging could not be attempted because pipe connection positions above the rig floor were unsuitable for breaking the string to insert logging tools. Efforts to sever the pipe were then delayed for 2.25 hr because eddy currents set off ASK system alerts. At 0225 hr on 29 June, the pipe was severed at 130.2 mbsf, between the top two joints of 5-in drill pipe. The pipe cleared the rig floor at 0500 hr, 7.75 hr after getting stuck, ending Hole 1071C.

Transit back to Site 1071
The ship was moved in dynamic positioning (DP) mode 1.9 nmi west from Site 1072, returning to Hole 1071C. A beacon (18.0 kHz) was deployed through the moonpool at 1910 hr on 7 July, followed by a second beacon (15.0 kHz) at 1940 hr.

Hole 1071D
In previous holes, Site 1071 was cored only with the APC/XCB system, and was plagued by unstable hole problems. However, the upper 350 mbsf at Site 1072 had been successfully cored with the smaller diameter (11-7/16 in. vs. 9-7/8 in.) rotary core barrel (RCB) coring system. We returned to Site 1071 in an attempt to RCB core down through the upper 250 mbsf and deepen the hole. The seafloor was tagged at 88 mbsl (99.4 mbrf), and Hole 1071D was spudded at 2100 hr on 7 July. RCB Cores 1072D-1R through 12R were cut from 0 to 85.7 mbsf; however, recovery was negligible despite taking half-cores. In an effort to improve recovery, the pump rate was reduced by 50%, but the pipe stuck going back to bottom with core barrel 13R. The pipe was worked with up to 150,000 lb overpull and 1100 amps (maximum allowed) without success, and the aft counter-balance sling on the top drive was broken in the process. The core barrel was removed through the top of the blocks, and the pipe was worked in tight hole about 6 m higher. The pipe came free while circulating at 1000 gpm and slacking off weight (until the pipe stood up in the elevators) with 1100 amps torque. The bit was pulled through the rotary at 1240 hr on 8 July.

Hole 1071E
A last attempt was made to drill down through the unstable sands at Site 1071 to log and core the remaining section quickly before hole instability became a problem. The ship was moved 15 m at 125° past hole 1071A, but was moved back 12 m to maintain positioning using two beacons. The seafloor was tagged at 87.6 mbsl (99.0 mbrf), and Hole 1071E was spudded at 1510 hr on 8 July, drilling to 259.8 mbsf using 20-bbl mud sweeps every 19 m. A short trip was made to 45 mbsf and found 15.6 m of soft fill. The hole was displaced with sepiolite mud, and the drilling BHA was pulled to change to a logging BHA rather than drop the bit in the hole. The bit change took 4 hr, and a VIT-TV survey of the seafloor showed a 3-m diameter crater with the beacon container from Hole 1071A in view and still upright. The logging bit was positioned at 45.8 mbsf after reaming fill at 39 m. The digital dual induction tool/long-spaced sonic/natural gamma spectrometry (DITE/LSS/NGTC) logging tool was unable to pass 60 mbsf (15 m below the bit). The logging bit was washed to 74 mbsf, and another attempt was made to log. The logging tool only got 7 m out of the pipe while the pipe was standing up in the elevators, indicating that it was stuck, with a heavy flow-back of seawater and fine sand indicating the hole was collapsing. The pipe was freed and pulled out, clearing the rotary at 2015 hr on 9 July, ending Hole 1071E.

Hole 1071F
Approval was received to move to two alternate sites in the MAT-8B hazards survey grid, which were projected to have higher silt and clay content in deeper objectives than proposed Site MAT-8B-3. Approval also was given to keep the same Site 1071 designation and drill through the unstable upper section to 250 m. The ship was moved in DP mode about 1 km northeast to proposed Site MAT-8B-4. A bottom survey was conducted over a flat, featureless seafloor.

The seafloor was tagged at 90 mbsl (101.5 mbrf), and Hole 1071F was spudded at 0027 hr on 10 July, drilling from 0 to 230.0 mbsf, where the pipe became stuck. The drill pipe was worked by setting weight down on the BHA, and using high torque (900 amps) and high circulation rates (700 gpm at 175 psi). The hole was reamed to 53.5 mbsf with the top drive, and 12 m of fill was cleaned out. Drilling continued from 230.0 to 233.5 mbsf, where the pipe became stuck again. The pipe was freed as before using high circulation rates and torques. Shortly thereafter, ASK system alarms shut down operations for 2.24 hr, after which the hole was then drilled ahead from 233.5 to 252.0 mbsf. RCB Cores 1071F-1R through 3R were recovered from 252.0 to 280.1 mbsf; however, torque and drag were high through this presumed sandy interval. The pipe was conditioned by reaming out and back in with the top drive, then drilling from 280.1 to 284.9 mbsf. Coring resumed with Core-4R (284.9-291.9 mbsf), but with no recovery. Therefore, the interval from 291.9 to 310 mbsf was drilled after which coring resumed with Cores 1071F-5R through 16R from 310.8 through 424.2 mbsf. Low recovery in Cores 12R to 16R suggested another interval of unconsolidated sands, so we decided it was prudent to cease coring operations and attempt to obtain logs to total depth.

A wiper trip was made to 43 mbrf, experiencing no drag or overpull with 48 m of fill. The bit was released in the hole, the hole was displaced with sepiolite mud, and the pipe was pulled out with the top drive to 53 mbsf for logging. During the first logging run, the LSS/DITE logging tool stopped 21 m below the end of pipe. The open-ended pipe was washed down to 87.7 mbsf; however, the hole appeared to be closing in rapidly. Therefore, logging was terminated and the pipe cleared the rotary at 0840 hr on 12 July ending Hole 1071F.

Hole 1071G
A final attempt was made to obtain logs using the LWD tools. The ship was moved 15 m at 35°. The seafloor was tagged at 90 mbsl (101.5 mbrf), and Hole 1071G was spudded at 1247 hr on 12 July. The hole was drilled with the LWD assembly to 95.1 mbsf in 4.5 hr; however, the pipe became stuck after making the next connection. Attempts to free the pipe with increased circulation rates, mud sweeps, and use of the jars were unsuccessful; therefore, it was assumed that the top of the hole had collapsed. Consequently, it was necessary to run the LWD Linc tool to retrieve the radioactive sources and download the logging data. Preparations were being made to sever the pipe, when the pipe came free and was retrieved. The bit cleared the rotary at 2315 hr, ending Hole 1071G.

To 174A Operations - Site 1072

To 174A Table of Contents

ODP Publications

ODP Homepage